OpenStudy (anonymous):

A soup recipe uses 3 pounds of meat to make 20 servings. How much is meet needed to make 15 servings?

OpenStudy (studygurl14):

Use proportions to solve. "3 pounds of meat to make 20 servings" aka, 3 pounds of meat PER 20 servings Per, in math language, means divide \(\large\rightarrow\Large\frac{3}{20}\) "How much is meet needed to make 15 servings?" aka, how much meat PER 15 servings \(\large\rightarrow\Large\frac{m}{15}\) Now, the ratios have to be equal, so... \(\large\rightarrow\Large\frac{3}{20}=\frac{m}{15}\) Solve for m. Do you know how to cross-multiply?